# Quantitative calibration of a TWPA applied to an optomechanical platform

### v1

@@ Abstract
In the last decade, the microwave quantum electronics toolbox has been enriched with quantum-limited detection devices such as Traveling Wave Parametric Amplifiers (TWPAs). The extreme sensitivity they provide is not only mandatory for some physics applications within quantum information processing, but is also the key element that will determine the detection limit of quantum sensing setups. In the framework of microwave optomechanical systems, an unprecedented range of small motions and forces is accessible, for which a specific quantitative calibration becomes necessary. We report on near quantum-limited measurements performed with an aluminum drumhead mechanical device within the temperature range 444~{}4mK -- 400400400~{}400mK. The whole setup is carefully calibrated, especially taking into account the power-dependence of microwave absorption in the superconducting optomechanical cavity. This effect is commonly attributed to Two-Level-Systems (TLSs) present in the metal oxide. We demonstrate that a similar feature exists in the TWPA, and can be phenomenologically fit with adapted expressions. The power and temperature dependence is studied over the full parameter range, leading to an absolute definition of phonon population (i.e. Brownian motion amplitude), with an uncertainty ±$\pm 20~{}\%$.
@@ I Introduction
For these usages, the TWPA detection enables the best possible discrimination between quantum bit states $|0\!>$ and $|1\!>$. In this respect, the read-out is essentially relative. But specific quantum sensing issues, especially linked to fundamental research, require an absolute quantitative calibration of the signal level coming out from the quantum amplifier. In 2024, the project Quantum Technologies for Neutrino Mass developed a sensing apparatus to measure the neutrino mass with a sensitivity of 10 $meV/c^{2}$ [18]. Improving this sensitivity is here a key ingredient and relies on quantum amplifiers. As well, one candidate for Dark Matter is the axion, which is today actively sought in cryogenic microwave setups employing TWPA read-out [19].
@@
In the present article we report on the quantitative characterization of a TWPA-based near quantum-limited detection setup. This is performed in the framework of a microwave opto-mechanical platform, able to cool down the system to the coldest possible temperatures achievable for condensed matter (by means of nuclear demagnetization [27]). The opto-mechanical measurements are performed using a drum-shaped aluminum membrane coupled to a superconducting cavity. The TWPA properties are analyzed on the basis of the Two Level Systems (TLS) theory [28, 29], phenomenologically reproducing results known for superconducting cavities [30]. It turns out that TLSs seem to play an important role in the dynamics of TWPAs, as recently shown with echo experiments [31]. We demonstrate that we can extract the population of the mechanical mode within ±$\pm 20~{}\%$ by means of an absolute calibration, reaching a population of 5 phonons at 4 mK.
@@ II Experimental setup
The microwave experimental platform is installed on the dilution unit of a nuclear demagnetization cryostat [27]. The schematics is reproduced in Fig. 1; it is the same setup as the one of Ref. [32]. A pump signal (which is required for the optomechanical schemes discussed in the next Section) is sent into the cryostat through attenuators that eliminate the thermal noise. Whenever an extra small probe tone is required for the experiment, it is injected through the same injection line. A double circulator is added in order to protect the sample cell from any signals traveling backwards from the amplifiers (noise, or reflected waves due to mismatched components). A second signal, created by splitting the pump tone at room temperature and amplitude-and-phase controlled, is injected in a separate line towards the amplifiers in order to suppress the strong pump. The two signals are combined by a directional coupler. Efficient pump cancellation is especially important here, since the TWPA saturates at extremely low input powers, typically about $-110~{}$dBm [33].
Figure 1: Low temperature microwave circuitry: after splitting the optomechanical pump signal into injection line and cancellation line, both signals are sent into the cryostat through attenuators and circulators. Both are recombined before the TWPA to avoid any saturation. Another circulator is implemented downstream to protect the TWPA from the noise generated by the cryogenic HEMT. The total attenuation in injection and gain in detection are calibrated within an uncertainty of about ±$\pm 1$ dB. The different anchoring stages of the dilution unit are specified.
The detection is performed by means of a Zurich Instruments® lock-in amplifier. The room temperature HEMT output signal is down-converted with a mixer and a Local Oscillator (LO) [27, 32]. The detection line that brings the signal to the lock-in is composed of a room temparature HEMT (a conventional Low Noise Factory® device), and a last generation Low Noise Factory® cryogenic HEMT with $T_{noise}\approx 2~{}$K. The first stage of amplification is performed with an ”Argo” Silent Waves® TWPA, with an insertion loss of about $3-4~{}$dB around 555~{}5GHz. This component enables us to bring the detection limit as close as possible to the Standard Quantum Limit (SQL). It requires a pump tone to function, which is sent through the cancellation line (Fig. 1). It also must be protected against the noise of the HEMT, which is why an extra circulator is present on the schematics between the two cryogenic amplifiers.
@@ III Basics of optomechanics
Figure 2: Typical optomechanical data taken at $T=20~{}$mK. a) Cavity measurement using a large probe tone ($1\cdot 10^{-4}~{}$nW, and no pump); the resonance is fit with a Lorentzian function defined in Eq. (1); note the logarithmic scale. b) Raw mechanical peak obtained with the blue pumping scheme ($P_{in}=2.5\cdot 10^{-3}~{}$nW), and its Lorentz fit. Integrating this peak leads to the sideband area $A_{sdb}$ expressed in photons/s, and described by Eq. (5), taking into account the proper gains and conversion factors. c) $\Gamma_{e\!f\!f}$ measurement using blue and red schemes (respectively anti-damping/damping processes) as a function of the on-chip applied power (corrected from injection attenuation). The absolute value of the slope directly gives $g_{0}$, the optomechanical coupling constant, through Eq. (4). d) Measured signal area for blue- and red- pumping configurations, with fits corresponding to Eq. (5). In c) and d) the dashed vertical corresponds to the threshold for self-oscillation (see text).
@@
with the sign $\pm$ standing for the two pump schemes (and $\hbar$ is Planck’s reduced constant). For blue pumping (-- sign), the damping goes down until it reaches zero: at that point, the mechanical mode starts to self-oscillate [39, 40]. The area of the mechanical sideband peak $A_{sdb}$ (in photons/s on chip) is given by [27]:
@@
with $G_{opt}$ the optomechanical gain function, and $A_{0}$ the original peak area in the absence of amplification/deamplification. This quantity can be converted from photons/seconds into an area $A_{ph}$ expressed in phonons:
@@ IV Cavity properties
Figure 3: Main: Total damping rate of the cavity measured as a function of pump applied power (blue scheme, using a small probe tone, $5\cdot 10^{-7}~{}$nW) at $T=150$ mK. The fit corresponds to Eq. (9). The gray vertical corresponds to the inflection point defined by the critical power $P_{cav}^{0}$. A dramatic drop ($\sim 50$) in the total damping rate is observed while increasing the power (see text). Inset: temperature dependence demonstrating the characteristic $\tanh$ Eq. (10) shape (see text).
@@
[…]
@@
Figure 4: Main: background noise measurements as a function of cryogenic temperature. The $T_{cryo}\rightarrow 0~{}$K interpolation corresponds to the fundamental limit of the detection system, here about 3.53.53.5~{}3.5SQL. Inset : Typical result of a TWPA pump parameter tuning scan at $T=20$ mK. The yellow part of the graph corresponds to the highest values of TWPA gain, while the red cross stands for the optimal set of parameters chosen (see text for further details), giving a gain of $18\pm 1~{}$dB.
Below about 300 mK, the $\kappa_{BCS}$ term is negligible. At our highest temperatures (about 400400400~{}400mK), the total damping rate $\kappa_{tot}$ reaches $\sim 3~{}$MHz. At the lowest temperatures, the TLS saturation can lead to about $50~{}\%$ change in cavity damping, which is significant and obviously needs to be considered for our quantitative approach. Besides, we show in the next Section that the same phenomenon happens in the TWPA: and a phenomenological fitting adapted from Ref. [30] describes well the experimental results.
@@ V Twpa properties
In Fig. 4 (main graph) we thus also plot the amplification chain total noise, referenced at the input of the TWPA. The graph is obtained from a careful calibration of the losses of all components inserted in the circuit, and the gains of the two HEMTs. The extracted noise figure had been validated against the cold/hot load technique; for details on the procedure, see Ref. [43]. We reach at the lowest temperatures about 3.5  photons system noise at 5 GHz, which is about the best figure for this amplifier.
@@
Figure 5: Main: TWPA transmission around 5 GHz (with TWPA pump off) measured as a function of applied probe power, for various temperatures (and optomechanics pump off). Fitting functions correspond to Eqs. (12,13). The drop reaches about $40~{}\%$ at 202020~{}20mK for low powers (see text). Inset: TWPA transmission at 150 mK (with TWPA pump off) measured with constant small probe power (about $3\cdot 10^{-17}~{}$), but as a function of the detuning of a large pump tone from the probe (optomechanics pump, about $2\cdot 10^{-12}~{}$). While detuning the pump, the losses reappear (for details, see text).
@@
Figure 6: Comparison of TWPA transmission versus power with/without TWPA pump at $T=150~{}$mK. Orange points correspond to Fig. 5 (TWPA off), while green points are measured turning on the TWPA pump (and scaling properly for the TWPA gain). The vertical band illustrates the saturation zone of the TWPA (see text for details).
Finally, Fig. 6 compares the transmission while turning on/off the TWPA pump. The two datasets start at the same level, by definition (this is actually how the gain is defined). But as the probe power is increased, we see a clear difference: when the TWPA pump is on, the signal very rapidly degrades. This is marked by the shaded region in Fig. 6: a first kink is visible around $10^{-16}~{}$, and then the signal drops dramatically above $10^{-14}~{}$. This is nothing but the TWPA amplifier’s saturation [33, 34]. What Fig. 6 tells us is that it is impossible to saturate the TLSs while keeping the TWPA gain optimum: the TWPA saturates first. As a result, if one wants a quantitative measurement of the transmitted signal amplitude, one has to take the TWPA TLSs into account: there is no way around here. To do so, phenomenological fits such as Eqs. (12,13) are enough, and will be applied in the next Section in the framework of optomechanics.
@@ VI Calibrated optomechanics signals
We now turn to the use of this microwave platform for optomechanics measurements. From Eq. (7), we extract the area $A_{ph}$. This is properly performed thanks to the cavity TLS analysis of Section IV, for various input powers and temperatures. At temperatures $T>1~{}$mK, blue and red pumping schemes are essentially equivalent and the value should match the Bose-Einstein phononic population:
@@
which reduces to $k_{B}T/(\hbar\,\Omega_{m})$ for $n_{ph}\gg 1$. We therefore plot in Fig. 7 the ratio $A_{ph}/n_{ph}$ as a function of cryostat temperature $T_{cryo}$. Note that the signal power is much smaller than $10^{-16}~{}$, and thus cannot saturate neither the TWPA, nor the TLSs (See Fig. 6). So without any correction, this should simply reflect the temperature dependence of the TWPA transmission $\delta$, Eqs. (12,13). To prove the point, we compare this to the same ratio obtained in a run with no TWPA, on the same cryostat [35]. The wiring was exactly the same, apart from the HEMT which was an old Caltech® model with much higher noise level (but essentially the same gain). We also plot the probe data of the previous Section.
@@
Finally, it is clear on Fig. 7 that the data reproducibility is about ±$\pm 20~{}\%$. This is actually intrinsic to the mechanical device, and arises from $1/f$-type fluctuations that limit the quality of the data averaging [35]. Such noises, which ultimately limit our relative error, deserve further studies. Relying on the work presented here, we can calibrate the setup over two decades in temperature, for any microwave power (and optomechanical scheme, as defined from the pump detuning). We end up at 444~{}4mK with an inferred population of about 5 phonons, reproducibly measured within ±$\pm 20~{}\%$, consistent with Eq. (14).
@@ VII Conclusions
This TLS dissipative channel appears as an extra insertion loss for the TWPA, which is power and temperature dependent. It is fit in order to produce a quantitative calibration of the detection line, in the full parameter space: one just needs to normalize the measured signal (for us, the optomechanical sideband area) with Eqs. (12,13). We demonstrate that the TWPA pump does not saturate the TLSs (the TWPA itself saturates first, and becomes unusable): the normalization procedure is therefore the only way to take the TLS effect into account. As such, we can extract from the measurements the mean phonon population of our nano-mechanical device from 444~{}4mK to 400400400~{}400mK, with an achieved uncertainty of ±$\pm 20~{}\%$ over the full temperature range.

### v2

@@ Abstract
In the last decade, the microwave quantum electronics toolbox has been enriched with quantum-limited detection devices such as Traveling Wave Parametric Amplifiers (TWPAs). The extreme sensitivity they provide is not only mandatory for some physics applications within quantum information processing, but is also the key element that will determine the detection limit of quantum sensing setups. In the framework of microwave optomechanical systems, an unprecedented range of small motions and forces is accessible, for which a specific quantitative calibration becomes necessary. We report on near quantum-limited measurements performed with an aluminum drumhead mechanical device within the temperature range 44~mK - 400400~mK. The whole setup is carefully calibrated, especially taking into account the power-dependence of microwave absorption in the superconducting optomechanical cavity. This effect is commonly attributed to Two-Level-Systems (TLSs) present in the metal oxide. We demonstrate that a similar feature exists in the TWPA, and can be phenomenologically fit with adapted expressions. If not taken into account, the error on the signal strength can be as large as a factor of about 2, which is unacceptable for quantitative experiments. The power and temperature dependence is studied over the full parameter range, leading to an absolute definition of phonon population (i.e. Brownian motion amplitude), with an uncertainty $\pm 20~\%$ limited by sources of noise internal to the optomechanical element.
@@ I Introduction
For these usages, the TWPA detection enables the best possible discrimination between quantum bit states |0>|0\!> and |1>|1\!>. In this respect, the read-out is essentially relative. But specific quantum sensing issues, especially linked to fundamental research, require an absolute quantitative calibration of the signal level coming out from the quantum amplifier. In 2024, the project Quantum Technologies for Neutrino Mass developed a sensing apparatus to measure the neutrino mass with a sensitivity of 10 $meV/c^{2}$ [18]. Improving this sensitivity is here a key ingredient and relies on quantum amplifiers. As well, one candidate for Dark Matter is the axion, which is today actively sought in cryogenic microwave setups employing TWPA read-out [19].
@@
[…]
@@ II Experimental setup
The microwave experimental platform is installed on the dilution unit of a nuclear demagnetization cryostat [27]. The schematics is reproduced in Fig. 1; it is the same setup as the one of Ref. [32]. A pump signal (which is required for the optomechanical schemes discussed in the next Section) is sent into the cryostat through attenuators that eliminate the thermal noise. Whenever an extra small probe tone is required for the experiment, it is injected through the same injection line. A double circulator is added in order to protect the sample cell from any signals traveling backwards from the amplifiers (noise, or reflected waves due to mismatched components). A second signal, created by splitting the pump tone at room temperature and amplitude-and-phase controlled, is injected in a separate line towards the amplifiers in order to suppress the strong pump. The two signals are combined by a directional coupler. Efficient pump cancellation is especially important here, since the TWPA saturates at extremely low input powers, typically about 110-110~dBm [33].
Figure 1: Low temperature microwave circuitry: after splitting the optomechanical pump signal into injection line and cancellation line, both signals are sent into the cryostat through attenuators and circulators. Both are recombined before the TWPA to avoid any saturation. Another circulator is implemented downstream to protect the TWPA from the noise generated by the cryogenic HEMT. The total attenuation in injection and gain in detection are calibrated within an uncertainty of about $\pm 1$ dB. The different anchoring stages of the dilution unit are specified.
The detection is performed by means of a Zurich Instruments® lock-in amplifier. The room temperature HEMT output signal is down-converted with a mixer and a Local Oscillator (LO) [27, 32]. The detection line that brings the signal to the lock-in is composed of a room temparature HEMT (a conventional Low Noise Factory® device), and a last generation Low Noise Factory® cryogenic HEMT with $T_{noise}\approx 2$~K. The first stage of amplification is performed with an ”Argo” Silent Waves® TWPA, with an insertion loss of about 343-4~dB around 55~GHz. This component enables us to bring the detection limit as close as possible to the Standard Quantum Limit (SQL). It requires a pump tone to function, which is sent through the cancellation line (Fig. 1). It also must be protected against the noise of the HEMT, which is why an extra circulator is present on the schematics between the two cryogenic amplifiers.
@@ III Basics of optomechanics
Figure 2: Typical optomechanical data taken at T=20T=20~mK. a) Cavity measurement using a large probe tone ($1\cdot 10^{-4}$~nW, and no pump); the resonance is fit with a Lorentzian function defined in Eq. (1); note the logarithmic scale. b) Raw mechanical peak obtained with the blue pumping scheme ($P_{in}=2.5\cdot 10^{-3}$~nW), and its Lorentz fit. Integrating this peak leads to the sideband area $A_{sdb}$ expressed in photons/s, and described by Eq. (5), taking into account the proper gains and conversion factors. c) $\Gamma_{e\!f\!f}$ measurement using blue and red schemes (respectively anti-damping/damping processes) as a function of the on-chip applied power (corrected from injection attenuation). The absolute value of the slope directly gives $g_{0}$, the optomechanical coupling constant, through Eq. (4). d) Measured signal area for blue- and red- pumping configurations, with fits corresponding to Eq. (5). In c) and d) the dashed vertical corresponds to the threshold for self-oscillation (see text).
@@
with the sign $\pm$ standing for the two pump schemes (and $\hbar$ is Planck’s reduced constant). For blue pumping (- sign), the damping goes down until it reaches zero: at that point, the mechanical mode starts to self-oscillate [39, 40]. The area of the mechanical sideband peak $A_{sdb}$ (in photons/s on chip) is given by [27]:
@@
with $G_{opt}$ the optomechanical gain function, and $A_{0}$ the original peak area in the absence of amplification/deamplification. This quantity can be converted from photons/seconds into an area $A_{ph}$ expressed in phonons, which corresponds to the occupation number of the mechanical mode:
@@ IV Cavity properties
Figure 3: Main: Total damping rate of the cavity measured as a function of pump applied power (blue scheme, using a small probe tone, $5\cdot 10^{-7}$~nW) at $T=150$ mK. The fit corresponds to Eq. (9). The gray vertical corresponds to the inflection point defined by the critical power $P_{cav}^{0}$. A dramatic drop ($\sim 50$ %) in the total damping rate is observed while increasing the power (see text). Inset: temperature dependence demonstrating the characteristic $\tanh$ Eq. (10) shape (see text).
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\kappa_{in}(T,P_{in})$ $\!\!\,=\!\!$ $\kappa_{TLS}(T)\frac{P_{cav}^{0}(T)/P_{in}}{1+P_{cav}^{0}(T)/P_{in}}+\kappa_{BCS}(T),$ (9)
$\!\!\!\kappa_{TLS}(T)$ $\,\,\!=\!\,$ $\kappa_{TLS}^{0}\tanh\left(\frac{\hbar\omega_{c}}{2k_{B}T}\right).$ (10)
@@
[…]
Below about 300 mK, the $\kappa_{BCS}$ term is negligible. At our highest temperatures (about 400400~mK), the total damping rate $\kappa_{tot}$ reaches $\sim 3$~MHz. At the lowest temperatures, the TLS saturation can lead to about $50~\%$ change in cavity damping, which is significant and obviously needs to be considered for our quantitative approach. Besides, we show in the next Section that the same phenomenon happens in the TWPA: and a phenomenological fitting adapted from Ref. [30] describes well the experimental results.
@@ V Twpa properties
[…]
@@
Figure 5: Main: TWPA transmission around 5 GHz (with TWPA pump off) measured as a function of applied probe power, for various temperatures (and optomechanics pump off). Fitting functions correspond to Eqs. (12,13). The drop reaches about $40~\%$ at 2020~mK for low powers (see text). Inset: TWPA transmission at 150 mK (with TWPA pump off) measured with constant small probe power (about $3\cdot 10^{-17}$~W), but as a function of the detuning of a large pump tone from the probe (optomechanics pump, about $2\cdot 10^{-12}$~W). While detuning the pump, the losses reappear (for details, see text).
@@
Figure 6: Comparison of TWPA transmission versus power with/without TWPA pump at T=150T=150~mK. Orange points correspond to Fig. 5 (TWPA off), while green points are measured turning on the TWPA pump (and scaling properly for the TWPA gain). The vertical band illustrates the saturation zone of the TWPA (see text for details).
Finally, Fig. 6 compares the transmission while turning on/off the TWPA pump. The two datasets start at the same level, by definition (this is actually how the gain is defined). But as the probe power is increased, we see a clear difference: when the TWPA pump is on, the signal very rapidly degrades. This is marked by the shaded region in Fig. 6: a first kink is visible around $10^{-16}$~W, and then the signal drops dramatically above $10^{-14}$~W. This is nothing but the TWPA amplifier’s saturation [33, 34]. What Fig. 6 tells us is that it is impossible to saturate the TLSs while keeping the TWPA gain optimum: the TWPA saturates first. As a result, if one wants a quantitative measurement of the transmitted signal amplitude, one has to take the TWPA TLSs into account: there is no way around here. To do so, phenomenological fits such as Eqs. (12,13) are enough, and will be applied in the next Section in the framework of optomechanics.
@@ VI Calibrated optomechanics signals
We now turn to the use of this microwave platform for optomechanics measurements. From Eq. (7), we extract the area $A_{ph}$. This is properly performed thanks to the cavity TLS analysis of Section IV, for various input powers and temperatures. At temperatures T>1T>1~mK, blue and red pumping schemes are essentially equivalent and the value should match the Bose-Einstein phononic population:
@@
which reduces to $k_{B}T/(\hbar\,\Omega_{m})$ for $n_{ph}\gg 1$. We therefore plot in Fig. 7 the ratio $A_{ph}/n_{ph}$ (measured phonon population normalized to expectation) as a function of cryostat temperature $T_{cryo}$. Note that the signal power is much smaller than $10^{-16}$~W, and thus cannot saturate neither the TWPA, nor the TLSs (See Fig. 6). So without any correction, this should simply reflect the temperature dependence of the TWPA transmission $\delta$, Eqs. (12,13). To prove the point, we compare this to the same ratio obtained in a run with no TWPA, on the same cryostat [35]: in this case, $A_{ph}/n_{ph}$ remains around 1 at all temperatures. The wiring was exactly the same, apart from the HEMT which was an old Caltech® model with much higher noise level (but essentially the same gain). We also plot the probe data of the previous Section.
@@
Finally, it is clear on Fig. 7 that the data reproducibility is about $\pm 20~\%$. This is actually intrinsic to the mechanical device, and arises from $1/f$-type fluctuations that limit the quality of the data averaging [35]. Such noises, which ultimately limit our relative error, deserve further studies. Relying on the work presented here, we can calibrate the setup over two decades in temperature, for any microwave power (and optomechanical scheme, as defined from the pump detuning). We end up at 44~mK with an inferred population of about 5 phonons, reproducibly measured within $\pm 20~\%$, consistent with Eq. (14).
@@ VII Conclusions
This TLS dissipative channel appears as an extra insertion loss for the TWPA, which is power and temperature dependent. It is fit in order to produce a quantitative calibration of the detection line, in the full parameter space: one just needs to normalize the measured signal (for us, the optomechanical sideband area) with Eqs. (12,13). We demonstrate that the TWPA pump does not saturate the TLSs (the TWPA itself saturates first, and becomes unusable): the normalization procedure is therefore the only way to take the TLS effect into account, which can lead to a loss in signal of up to a factor 2 here. As such, we can extract from the measurements the mean phonon population of our nano-mechanical device from 44~mK to 400400~mK, with an achieved uncertainty of $\pm 20~\%$ over the full temperature range. This scatter is actually neither related to the TWPA nor the method we employ, it is actually internal to the optomechanical device [35].